Two balls of identical mass of 0.502 kg and with identical charges placed on them hang...

Charged Balls

Two balls of identical mass of 0.502 kg and with identical charges placed on them hang from the ceiling on identical strings of lengths 0.710 m as shown. If the angle with respect to the vertical that the strings form is 29.0 degrees, what is the charge on each ball? Please give your answer in micro-coulomb (i.e. do not enter a unit, but calculate your answer in micro-coulomb).
Answer #1

During equilibrium, equating forces in vertical direction:

T*cos(θ) = mg ---- (1)

where T = tension in the string

Similarly, equating forces in the horizontal direction:

T*sin(θ) = E*q = (k*q/d^2)*q = k*q^2/d^2 (electrostatic force)

Also, d/2 = L*sin(θ)

So, d = 2Lsin(θ)

So, dividing equation 1 from equation 2:

tan(θ) = Eq/mg = ( kq^2/(2L*sinθ)^2) / (mg)

=> tan (29 deg)= (9*10^9*q^2/(2*0.71*sin(29 deg))^2) / (0.502*9.8)

So, q = 1.1983*10^-5 C = 11.983 uC

Answer = 11.983